notice a water ballon fall past your classroom window. You estimate that it took the balloon about t seconds to fall the length of the window and that the window is about y meters high. Suppose the balloon started from rest. Approximately how high above the top of the window was it released?
1
Expert's answer
2019-09-24T09:02:29-0400

Let's denote the distance between the top of the window and the start point of balloon by "h". We shall use the law for uniformly accelerated motion. Let's assume that the balloon travelled this distance "h" in some time "\tau" (we don't know it). Then it travelled distance "h + y" in time "{\tau + t}" (we know it from the conditions of the problem - it travelled the distance "y" in time "t" ). The law for uniformly accelerated motion give us two equation for these cases (we use that the balloon started from rest thus "{v_0} = 0" and let the initial point of the coordinate system at the initial position of the ballon)


"\left\{ \begin{array}{l}h = \frac{{g{\tau ^2}}}{2}\\h + y = \frac{{g{{(\tau + t)}^2}}}{2}\end{array} \right."

Using the first equation we get


"\tau = \sqrt {\frac{{2h}}{g}}"

Let's substitute this to the second equation and simplify it


"h + y = \frac{g}{2}{(\sqrt {\frac{{2h}}{g}} + t)^2} = \frac{g}{2}(\frac{{2h}}{g} + 2t\sqrt {\frac{{2h}}{g}} + {t^2}) = h + \frac{{g{t^2}}}{2} + t\sqrt {2hg}"

Transfer all terms except the last one from the right to the left side


"t\sqrt {2hg} = y - \frac{{g{t^2}}}{2}"

squaring

"2hg{t^2} = {(y - \frac{{g{t^2}}}{2})^2}"

and expressing "h"


"h = \frac{{{{(y - \frac{{g{t^2}}}{2})}^2}}}{{2g{t^2}}}"


This is the answer

P.S. It can be expressed in different (but equivalent) forms


"h = \frac{{{{(y - \frac{{g{t^2}}}{2})}^2}}}{{2g{t^2}}} = \frac{{{{(2y - g{t^2})}^2}}}{{8g{t^2}}} = \frac{{g{t^2}}}{8} - \frac{y}{2} + \frac{{{y^2}}}{{2g{t^2}}}"